# OmniVLA: Physically-Grounded Multimodal VLA with Unified Multi-Sensor Perception for Robotic Manipulation

### v2

@@ IV Evaluation
We evaluate OmniVLA with a real-world prototype across several sensor-related manipulation tasks. First, we show OmniVLA’s unique beyond-RGB perception capabilities, using daily tasks that require non-visible cues, including thermal-based pick-and-place, see-through and open the box with an object inside, and uncovering ringing items. OmniVLA significantly outperforms RGB-only VLA models and VLA models trained with unprocessed sensor images. We also highlight the data efficiency of our approach. Second, we show that our approach provides superior generalization capability for sensor-related tasks, outperforming baseline approaches.
@@
Figure 5: Examples of Robotic Manipulation Task Completion over Time. (a) Thermal: finding the cold drink. (b) mmWave: opening the box with object inside. (c) Acoustic: uncovering the ringing phone. The first three images are sensor-masked images; the rest images are raw RGB images for visibility.
@@
TABLE I: Success rates and task scores across three types of sensor-modality tasks.
@@ IV-A Experimental Setup
Model training and inference. We use SmolVLA [2] as the base model by default to implement our design and use the pre-trained weights. We expect our approach to be compatible with most existing RGB-only VLA models. We use multiple Nvidia A100 GPUs on a server for distributed training and use a local RTX 4090 GPU for model inference during system evaluation and inference. For real-time inference, we load both the VLA model and segmentation model on the local RTX 4090 machine, which is able to output 15 predictions per second for end-to-end action prediction. We expect to have lower delays with code implementation optimization. Note that the VLM model for the segmentation prompt is invoked only once per task to avoid action prediction delay.
@@
Thermal mmWave Acoustic Average Thermal mmWave Acoustic Average
@@
TABLE II: Model comparisons evaluated across three types of sensor tasks.
@@
TABLE III: Pretraining effectiveness by comparing per-stage success rates (%). Gains are absolute improvements over two baselines.
Evaluation metrics and baselines. We evaluate model performance using task success rates computed over 25 independent trials per task, complemented by task scores: 0.5 score for choosing the right item to interact with, 0.5 score for performing the correct manipulation, e.g., picking up and placing in the container, opening up the box, removing the coverings. For baselines, we compare our approach against (1) VLA-RGB: VLA models with standard RGB input only for training and inference, (2) VLA-RAW: VLA models with raw sensor data/images input for training and inference. It uses the same model architecture as OmniVLA, but skips the segmentation and overlay step for the sensor data processing. Note that we still apply beamforming for VLA-RAW model mmWave and acoustic sensor input to have a strong baseline.
@@ IV-B Multi-sensory Task Performance
We first evaluate OmniVLA performance on manipulation tasks compared with the baselines. Then we evaluate the performance over different foundational robotic models. Finally, we explore the impact of finetuning data amount on the final performance.
Overall Performance. We evaluate OmniVLA after training on 100 expert demonstration episodes of thermal and acoustic modality tasks and 200 episodes of mmWave modality task individually. mmWave task requires more training data because opening a box is more difficult to learn compared to typical pick-and-place type of actions, which seldom appears in current VLA model pretraining dataset. For each demonstration, we randomize objects’ positions on a table within the robot’s workspace. As SmolVLA pretraining dataset does not include any non-RGB sensor, we consider the number of episodes are reasonable and shows high learning efficiency of our approach. Table I shows success rates and task scores across tasks. OmniVLA finishes the tasks requiring extra sensors successfully at a high rate of 84% on average. OmniVLA consistently outperforms all baseline configurations across the three tasks, demonstrating the effectiveness of our unified multi-sensory perception approach. On average, we outperforms VLA-RGB model and VLA-RAW model by 59% and 28% in success rate, respectively. OmniVLA also improves the task score by 0.45 and 0.17, respectively. This shows sensor-modalities effectively provided new capabilities for the VLA model and our sensor-masked image representation significantly boosts performance.
Comparing base models. We apply our approach to Pi0 [1] and compare the performance across 3 tasks. Table II shows that both models work, while smolVLA provides better performance. This is likely due to smolVLA is pre-trained with lerobot robot arm dataset. Overall, this shows OmniVLA’s potential for combining with various current VLA models to provide extra new sensor-modality capabilities without starting from scratch.
@@
Learning efficiency. We quantitatively compare the learning efficiency of using sensor-masked images and raw sensor data (no overlaying on RGB images), by training with an increasing number of thermal task episodes. As shown in Figure 6, OmniVLA constantly outperforms VLA-RAW model, achieving a similar success rate with only around 50% of the training episodes. This shows the data efficiency of our proposed sensor-masked image representation, which significantly reduces the required finetuning data amount compared with raw sensor images.
@@ IV-C Generalization Performance
Figure 7: Success rates when adapting to unseen tasks. We compare the pretrained OmniVLA model with two baselines, VLA-Base: no pretraining, Pretrained VLA-Raw: pretrained with raw sensor images.
Multi-sensory pretraining. We construct a pretraining corpus of 800 demonstration episodes: 200 episodes for each of three sensor-modality tasks (a total of 600) and 200 episodes for generic pick-and-place with everyday objects.
We pre-train our OmniVLA model with the mixed dataset, then test the performance after performing few-shot learning with only 25 demonstration episodes of performing unseen tasks (matching common RGB-only VLA testing practice). We induce distribution shifts in object identity and materials on the basis of tasks shown in Figure 5): (1) swap the drink type, (2) replace cardboard with foam boxes, and (3) substitute T-shirts with towels. The object locations are randomized within the feasible workspace of the robot arm and towels are randomly folded.
Baselines. To show the effectiveness, we implement two baselines: (1) VLA-Base — no multi-sensory pretraining, using the same OmniVLA model architecture with RGB-only model (SmolVLA) as backbone weights, finetuned on the 25 demonstrations. (2) Pretrained VLA-RAW — OmniVLA model pretrained on the same 800 episodes but feed raw sensor heatmaps/images directly (no segmentation/overlay), then perform the same 25-shot adaptation. Our method, OmniVLA, uses the identical pretraining and few-shot protocols but with sensor-masked images as input. To compare different methods in more detail, we decompose each task into two stages: Stage 1 (select the correct target to interact with) and Stage 2 (complete the subsequent manipulation). We report stage-wise success rates and scores (Table III), and overall task successful rate (Figure 7).
As shown in Table III, OmniVLA shows substantially better generalization to unseen sensor-modality tasks than both baselines. In Stage 1, OmniVLA improves success rates by 17% over VLA-Base and by 32% over Pretrained VLA-RAW on average. In Stage 2, OmniVLA yields gains of up to 76%. These results suggest two complementary effects: (i) the sensor-masked overlay makes sensor cues spatially aligned with RGB, which simplifies learning the selection policy (Stage 1); and (ii) large-scale multi-sensory pretraining supplies transferable manipulation priors that boost few-shot control (Stage 2). Consequently, few-shot task success rate increases by 59% over VLA-Base and 28% over Pretrained VLA-RAW on average, and up to 68% across the three unseen tasks, as shown in Figure 7. To conclude, both sensor-masked images and pretraining significantly improve the success rate on unseen tasks with a few episodes of finetuning, showing powerful generalization capability.
@@ V Conclusion
We introduce OmniVLA, a multi-modal vision-language-action model that equips embodied AI systems with perception capabilities beyond the visible spectrum by integrating sensors, i.e., infrared, mmWave, and acoustic sensors. Our core design is the sensor-masked image, a unified representation that spatially grounds and semantically aligns diverse sensor data onto RGB images. This image-native approach allows reusing pretrained vision encoder, enabling data-efficient learning with lightweight per-sensor projection layers. Through extensive real-world experiments, we demonstrated that OmniVLA significantly outperforms baseline models on challenging manipulation tasks, showing an average success rate of 84%, exceeding the 25% of an RGB-only VLA and the 56% of a model trained on unprocessed sensor data. OmniVLA is a step towards creating more versatile and perceptive embodied AI that can fully understand and interact with their physical surroundings.

### v3

@@ IV Evaluation
We evaluate OmniVLA with a real-world prototype across several sensor-related manipulation tasks. First, we show our unique beyond-RGB perception capabilities, using daily tasks that require non-visible cues, including thermal-based pick-and-place, see-through and open the box containing the object, and uncovering hidden ringing items. OmniVLA significantly outperforms RGB-only VLA models and VLA models trained with unprocessed sensor images. We also highlight the data efficiency of our approach. Second, we show that our approach provides superior generalization capability for sensor-related tasks, outperforming baselines.
@@
Figure 5: Examples of Robotic Manipulation Task Completion over Time. (a) Thermal: finding the cold drink. (b) mmWave: opening the box with an object inside. (c) Acoustic: uncovering the ringing phone. The first three images are sensor-masked images; the rest of the images are raw RGB images for visibility.
@@
TABLE I: OmniVLA success rates and task scores across three types of sensor-modality tasks.
@@ IV-A Experimental Setup
Model training and inference. We use SmolVLA [34] as the base model by default to implement our design and use the pre-trained weights. We expect our approach to be compatible with most existing RGB-only VLA models. We use 8 NVIDIA A100 GPUs on a server for distributed training and use a local RTX 4090 GPU for model inference during system evaluation. Depending on the size of the demonstration dataset, the time required for training is around 14 hours for 50 K optimization steps with batch size 32. For real-time inference, we load both the VLA model and segmentation model on the local RTX 4090 machine, which is able to output 15 predictions per second for end-to-end action prediction. Note that prompting the VLM model for segmentation keywords is invoked at the beginning of the task and only updated infrequently to minimize overhead. We expect to have lower delays with code optimization.
@@
Base Model Thermal mmWave Acoustic Average Thermal mmWave Acoustic Average
@@
TABLE II: OmniVLA performance when using two different base models as backbone, evaluated across three types of sensor tasks.
@@
TABLE III: Pretraining effectiveness by comparing per-stage success rates (%). Gains are improvements over baselines.
Evaluation metrics and baselines. We evaluate model performance using task success rates computed over 25 independent trials per task with random object placement, complemented by task scores: 0.5 score for choosing the right item to interact with, 0.5 score for performing the correct manipulation, e.g., picking up and placing in the container, opening up the box, and removing the coverings. For baselines, we compare our approach against the following ablation baselines: (1) VLA-RGB (modality ablation): VLA models with standard RGB input only for training and inference, without our architecture changes. Since we use SmolVLA as the backbone, this baseline model is essentially a SmolVLA model fine-tuned on our task dataset. (2) VLA-RAW (representation ablation): VLA models with raw sensor data/images for training and inference input. It uses the same model architecture as OmniVLA, but skips the segmentation and overlay step for the sensor data processing. Note that we still apply beamforming for the VLA-RAW model mmWave and acoustic sensor input to have a strong baseline.
@@ IV-B Multi-sensory Task Performance
We first evaluate OmniVLA performance on manipulation tasks compared with the baselines. Then we evaluate the performance over different foundational robotic models. Finally, we explore the impact of finetuning data amount on final performance.
Overall Performance. We evaluate OmniVLA after training on 100 expert demonstration episodes of thermal and acoustic modality tasks and 200 episodes of mmWave modality task individually. The mmWave task requires more training data because opening a box is more difficult to learn compared to typical pick-and-place type of actions, which seldom appears in VLA model pretraining datasets. For each demonstration, we randomize objects’ positions on a table within the robot’s workspace. As the SmolVLA pretraining dataset does not include any non-RGB sensor, we consider the number of episodes reasonable and showing high learning efficiency of our approach. Table I shows success rates and task scores across tasks. OmniVLA finishes the tasks requiring extra sensors successfully at a high rate of 84% on average. OmniVLA consistently outperforms all baseline configurations across the three tasks, demonstrating the effectiveness of our unified multi-sensory perception approach. On average, OmniVLA outperforms VLA-RGB model and VLA-RAW model by 59% and 28% in success rate respectively. OmniVLA also improves the task score by 0.45 and 0.17, respectively. This shows that sensor-modalities effectively provide new capabilities for the VLA model and our sensor-masked image representation significantly boosts performance.
Comparing base models. To evaluate the applicability of our approach on different base VLA models, we apply it to Pi0 [3] and compare the performance across 3 tasks. Table II shows that both models work, while SmolVLA provides better performance. This is likely due to SmolVLA is pre-trained with lerobot robot arm dataset [34]. Overall, this shows OmniVLA’s potential for combining with various RGB-only VLA models to provide extra new sensor-modality capabilities without starting from scratch.
@@
Learning efficiency. We quantitatively compare the learning efficiency of using sensor-masked images and raw sensor data (no overlaying on RGB images), by training with an increasing number of thermal task episodes. As shown in Figure 6, OmniVLA constantly outperforms VLA-RAW model, achieving similar success rate with only around 50% of the training episodes. This shows the data efficiency of our proposed sensor-masked image representation, which significantly reduces the required finetuning data amount compared with using raw sensor images.
@@ IV-C Generalization Performance
Figure 7: Success rates when adapting to unseen tasks. We compare the pretrained OmniVLA model with two baselines, OmniVLA-Base (no pretraining), Pretrained VLA-RAW (pretrained with raw sensor images).
Multi-sensory pretraining. We construct a pretraining corpus of 800 demonstration episodes: 200 episodes for each of three sensor-modality tasks (total 600) and 200 episodes for generic pick-and-place with everyday objects.
We pre-train our OmniVLA model with the mixed dataset, then test the performance after performing few-shot learning with only 25 demonstration episodes of performing unseen tasks (matching common RGB-only VLA testing practice). We induce distribution shifts in object identity and materials based on tasks shown in Figure 5): (1) swap the drink type, (2) replace cardboard with foam boxes, and (3) substitute T-shirts with towels. The object locations are randomized within the feasible workspace of the robot arm and towels are randomly folded.
Baselines. To show the effectiveness, we implement two ablation baselines: (1) OmniVLA-Base — no multi-sensory pretraining, using OmniVLA model architecture with SmolVLA backbone weights, finetuned on the 25 demonstrations. (2) Pretrained VLA-RAW — OmniVLA model pretrained on the same 800 episodes but feed raw sensor heatmaps/images directly (no segmentation/overlay), then perform the same 25-shot adaptation. Our method, OmniVLA, uses the identical pretraining and few-shot protocols but with sensor-masked images as input. To compare different methods in more detail, we decompose each task into two stages: Stage 1 (select the correct target to interact with) and Stage 2 (complete the subsequent manipulation). We report stage-wise success rates and scores (Table III), and overall task success rate (Figure 7).
As shown in Table III, OmniVLA shows substantially better generalization to unseen sensor-modality tasks than both baselines. In Stage 1, OmniVLA improves success rates by 17% over OmniVLA-Base and by 32% over Pretrained VLA-RAW on average. In Stage 2, OmniVLA yields gains of up to 76%. These results suggest two complementary effects: (i) the sensor-masked overlay makes sensor cues spatially aligned with RGB, which simplifies learning the selection policy (Stage 1); and (ii) large-scale multi-sensory pretraining supplies transferable manipulation priors that boost few-shot control (Stage 2). Consequently, the few-shot task success rate increases by 59% over OmniVLA-Base and 28% over Pretrained VLA-RAW on average, and up to 68% across the three unseen tasks, as shown in Figure 7. To conclude, both sensor-masked images and pretraining significantly improve the success rate on unseen tasks with a few episodes of finetuning, showing powerful generalization capability.
@@ V Conclusion
We introduce OmniVLA, a multi-modal vision-language-action model that equips robots with perception capabilities beyond the visible spectrum by integrating sensors, i.e., infrared, mmWave, and acoustic sensors. Our core design is the sensor-masked image, a unified representation that spatially grounds and semantically aligns diverse sensor data onto RGB images. This image-native approach allows reusing a pretrained vision encoder, enabling data-efficient learning with lightweight per-sensor projection layers. Through extensive real-world experiments, we demonstrated that OmniVLA significantly outperforms baseline models on challenging manipulation tasks, showing an average success rate of 84%, exceeding the 25% of an RGB-only VLA and the 56% of a model trained on unprocessed sensor data. OmniVLA is a step towards creating more versatile and perceptive robots that can fully understand and interact with their physical surroundings.